Compute the extension in length of a wire 530 cm long and a diameter of 0.25 cm. When the body of mass 540 gm is applied vertically. The young modulus of given wire is 5x109 Dyne/cm2.


1
Expert's answer
2021-06-22T09:47:50-0400

The stress in the wire:


"\sigma=\frac FA=\frac{mg}{A}=\frac{4mg}{\pi d^2}."

Young's modulus in terms of stress and strain:


"E=\frac\sigma\epsilon,\\\space\\
\epsilon=\frac \sigma E=\frac{\Delta l}{l},\\\space\\
\Delta l=l\frac\sigma E."

Substitute the stress:


"\Delta l=l\frac{4mg}{\pi d^2E}=1.1\text{ cm}."

Do not forget to convert all units to SI!